If binomial distribution X~Bin(50,0.4) coincides with normal distribution X*~N(μ,σ2) , then find the value of μ, σ and

P[X ≤ 5] , P[X ≤ 15] , P[5 ≤ X ≤ 15] , P[5 < X < 15].


1
Expert's answer
2021-12-07T13:31:28-0500
"\mu=np=50(0.4)=20""\sigma^2=npq=50(0.4)(1-0.4)=12"





"\sigma=\sqrt{12}=2\sqrt{3}"

Continuity Correction Factor 

Using Normal Distribution with Continuity Correction "P(X\leq 5)"

Using Normal Distribution with Continuity Correction



"P(X<5.5)=P(Z<\dfrac{5.5-20}{2\sqrt{3}})""\approx P(Z<-4.18579)\approx0.00001421"

Using Normal Distribution with Continuity Correction "P(X\leq 15)"

Using Normal Distribution with Continuity Correction



"P(X<15.5)=P(Z<\dfrac{15.5-20}{2\sqrt{3}})""\approx P(Z<-1.299038)\approx0.096965"



Using Normal Distribution with Continuity Correction "P(5\leq X\leq 15)=P(X\leq 15)-P(X<5)"

Using Normal Distribution with Continuity Correction



"P(X<15.5)-P(X<4.5)""=P(Z<\dfrac{15.5-20}{2\sqrt{3}})-P(Z<\dfrac{4.5-20}{2\sqrt{3}})""\approx P(Z<-1.299038)-P(Z<-4.474465)""\approx0.096965-0.00000383""\approx0.096961"

Using Normal Distribution with Continuity Correction "P(5< X<15)=P(X< 15)-P(X\leq5)"

Using Normal Distribution with Continuity Correction



"P(X<14.5)-P(X<5.5)""=P(Z<\dfrac{14.5-20}{2\sqrt{3}})-P(Z<\dfrac{5.5-20}{2\sqrt{3}})""\approx P(Z<-1.587713)-P(Z<-4.18579)""\approx0.05617565-0.00001421""\approx0.05616144"